### v1

# Breaking the Performance Ceiling in Complex Reinforcement Learning requires Inference Strategies

Abstract
Reinforcement learning (RL) systems have countless applications, from energy-grid management to protein design. However, such real-world scenarios are often extremely difficult, combinatorial in nature, and require complex coordination between multiple agents. This level of complexity can cause even state-of-the-art RL systems, trained until convergence, to hit a performance ceiling which they are unable to break out of with zero-shot inference. Meanwhile, many digital or simulation-based applications allow for an inference phase that utilises a specific time and compute budget to explore multiple attempts before outputting a final solution. In this work, we show that such an inference phase employed at execution time, and the choice of a corresponding inference strategy, are key to breaking the performance ceiling observed in complex multi-agent RL problems. Our main result is striking: we can obtain up to a 126% and, on average, a 45% improvement over the previous state-of-the-art across 17 tasks, using only a couple seconds of extra wall-clock time during execution. We also demonstrate promising compute scaling properties, supported by over 60k experiments, making it the largest study on inference strategies for complex RL to date. Our experimental data and code are available at https://sites.google.com/view/inf-marl [^1] A GitHub repository will be released shortly..
Figure 1: Improvement from using inference-time search over zero-shot state-of-the-art. Across 17 complex reinforcement learning tasks, we obtain consistent and significant performance gains using only a 30 second search budget during execution.
1 Introduction
Learning to solve sequential decision-making tasks is a central challenge in artificial intelligence (AI), with far-reaching applications ranging from energy-grid optimisation (Ahmad et al., 2021) and autonomous logistics (Laterre et al., 2018) to molecular discovery (Olivecrona et al., 2017) and drug design (Popova et al., 2018). Complex sequential real-world problems that cannot be solved by traditional optimisation techniques are inherently complex and require navigating high-dimensional solution spaces. Reinforcement Learning (RL) presents a promising avenue to improve our capacity to find efficient solutions to these problems, but despite stunning progress over the past decade, such as human-level performance in Atari games (Mnih et al., 2015), defeating the world champion in the game of Go (Silver et al., 2016) or aligning AI systems with human preferences (Stiennon et al., 2020), current approaches are facing challenges that prevent their common deployment in most real-world systems (Dulac-Arnold et al., 2020).
A major source of this difficulty lies in the combinatorial nature of many decision-making tasks. As the problem size increases, the space of possible solutions grows exponentially (Karp, 1975). In multi-agent systems, the challenge compounds: agents must coordinate in environments where only partial information is available, the joint action space is combinatorial, and optimal behaviour depends on precise interaction with other agents (Bernstein et al., 2000; Canese et al., 2021). These properties make it fundamentally difficult to rely on the zero-shot performance of a trained policy, even if that policy was optimised to convergence on a representative training distribution. This causes the gap between zero-shot performance and optimality to grow substantially with increasing complexity (see Fig. 1).
However, numerous practical applications are not restricted to producing a single zero-shot solution. Instead, inference is often permitted to take place over a few seconds, minutes or hours, with a given computational resource. Furthermore, models and simulators are often accessible and very efficient (e.g., energy grid management, train scheduling, package delivery, routing, printed circuit board design) and provide either an exact score or a very accurate approximation. In other applications where the gap to reality may be larger (e.g., protein design, robotics), improving the solution under the simulated score can still arguably provide significant improvement towards the real objective (Hayes et al., 2025; Dona et al., 2024; Hundt et al., 2019; Rao et al., 2020).
This opens up an opportunity: rather than relying on a single attempt of the trained policy, the time budget and compute capacity can be leveraged to actively search for better solutions using multiple attempts, following an inference-time strategy. For instance, progressively building a tree of possible solutions, or adapting the policy using outcomes of past attempts. Even straightforward strategies can provide significant performance improvement with low time cost. For instance, generating a large batch of diversified solutions in parallel, using stochastic sampling, rather than a single greedy solution: given a modern GPU, this enables to produce hundreds of solutions, for the same wall-clock time, enabling massive exploration at no time cost. These strategies are rarely emphasized in existing benchmarks (Papoudakis et al., 2021; Mahjoub et al., 2025), and many practitioners invest months of research trying to improve the zero-shot performance of their models on scenarios where only marginal improvement may still be achieved. Whereas they could unlock performance gains from inference-time search, at negligible wall-clock time cost with moderate compute capacity (Fig. 1).
Research in RL for Combinatorial Optimisation (CO) has produced efficient inference strategies (Bello et al., 2016; Hottung et al., 2022; Choo et al., 2022; Chalumeau et al., 2023b), often referred to as active search, or online adaptation methods. However, their empirical study is still limited to a few problems, a narrow range of budget settings (Chalumeau et al., 2024b), and barely no insight on their scaling properties. In the multi-agent case, there is no study on inference strategies for collaborative teams: most work on team adaptation focus on Ad Hoc Teamwork (Mirsky et al., 2022; Wang et al., 2024a; Ruhdorfer et al., 2025), which is adjacent to our objective. Interestingly, most recent studies about the impact of inference strategies come from the Large Language Model (LLM) literature (Snell et al., 2025; Muennighoff et al., 2025; Wu et al., 2025), where the adequate combination of efficient models with inference strategies is currently state-of-the-art (SOTA).
In this work, we formalise and investigate the role of inference strategies in complex decision-making tasks. To capture the full complexity of tasks described above, we formulate our problem setting as a decentralised partially observable Markov decision process (Dec-POMDP) (Kaelbling et al., 1998). This is instead of the typical single-agent MDP used in many RL studies. We make this choice for several reasons: (1) it more accurately maps onto many complex real-world problems of interest, (2) Dec-POMDPs subsume MDPs by being strictly more complex (Bernstein et al., 2000), and (3) because of this, we expect our findings to translate to all simpler problem formulations. Within this setting, we provide a unifying view of popular inference-strategy paradigms, including policy sampling, tree search (Choo et al., 2022), online fine-tuning (Bello et al., 2016; Hottung et al., 2022), and diversity-based search (Chalumeau et al., 2023b). Strikingly, we show that across a wide range of specifically selected difficult RL problems, inference strategies boost performance on average by $45$%, over zero-shot SOTA. Furthermore, in the best of cases, this boost can be as large as $126$%. All of this, using only a couple of seconds of additional execution time.
Our results call for a shift in how RL systems are evaluated and deployed: inference strategies are not a minor post-processing step, but a key performance driver in realistic conditions. This work sets the foundation for a more nuanced view of inference in sequential decision-making and provides the tools to build systems that can scale with compute. All our code and experimental data can be accessed at: https://sites.google.com/view/inf-marl.
2 Related work
Inference Strategies from RL for CO
Beyond naive stochastic sampling, several paradigms have been explored to generate the best possible solution using a trained policy checkpoint during inference. Online fine-tuning Bello et al. (2016) retrains all policy parameters with RL using past attempts. Hottung et al. (2022) re-trains only a subset of the policy’s parameters to reduce memory and compute overheads, enabling more attempts for a given inference budget. It also adds a imitation learning term to the RL term, to force exploration close to the best solution found so far. Choo et al. (2022) uses tree search, with simulation guided node estimates under budget constraints, which outperforms Beam Search and Monte Carlo Tree Search (Coulom, 2006). Diversity-based methods: inspired by previous diversity-seeking approaches, like unsupervised skill discovery (Eysenbach et al., 2019; Sharma et al., 2019; Kumar et al., 2020) and quality-diversity (Chalumeau et al., 2023a; Cully and Demiris, 2017), Grinsztajn et al. (2023) introduces an RL objective that trains a population of diverse and specialized policies, efficient for few-shot performance. Chalumeau et al. (2023b) uses this objective and encodes the diversity in a continuous latent space that can be searched at inference-time, introducing the SOTA method compass; meanwhile Hottung et al. (2024) uses a similar approach but with a discrete encoding space.
These works have introduced most of the inference strategies we consider in this paper, but they fall short on three important aspects that we aim to improve: (i) they evaluate inference strategies on benchmarks where over $95$% zero-shot optimality is already achieved, leaving little room for meaningful gains, (ii) these benchmarks rely on domain-specific tricks such as starting points or instance augmentations; and (iii) methods are compared under a unique budget setting, overlooking the fact that relative performance depends on the available compute and time budget. In addition, their ability to scale with compute remains unexplored, despite being a critical property.
Search and adaptation in Multi-Agent RL
There is only limited work on inference strategies for MARL. Most work about search and adaptation within MARL focus on the challenge of Ad Hoc Teamwork (Yourdshahi et al., 2018; Hu et al., 2020; Mirsky et al., 2022; Wang et al., 2024a; Ruhdorfer et al., 2025; Hammond et al., 2025), often in the form of zero-shot coordination, where agents must generalize to new partners at execution time. While these lines of work share some methodological similarities, for instance using diversity-seeking training (Long et al., 2024; Lupu et al., 2021) or adapting through tree search (Yourdshahi et al., 2018), they pursue fundamentally different goals and remain orthogonal. In our work, our focus is primarily on solving difficult and complex industrial optimisation tasks.
Inference-time compute for LLMs
Recent advances in LLMs are closely intertwined with the use of inference strategies (Snell et al., 2025; Wei et al., 2022; Wang et al., 2024b), and a growing effort has gone into studying their scaling properties (Muennighoff et al., 2025; Wu et al., 2025). However, the typical inference-time setting is usually different from ours. LLMs have very costly forward passes, and cannot access the exact score of their answers, but can approximate them using a reward model (Ouyang et al., 2022). Most popular strategies for LLMs are designed for few shots, namely sampling and ensembling (e.g., majority voting).
Overall, numerous inference strategies have been proposed in the literature, yet their efficiency under various evaluation settings remains unexamined. Multi-agent RL, despite its inherent complexity, rarely considers inference-time search beyond Ad Hoc Teamwork. Moreover, the broader field of decision-making has not systematically studied how inference strategies scale with compute. Our work aims at filling these gaps by extending the evaluation of inference strategies in RL, demonstrating major performance gains over a wide range of budget settings with impressive scaling properties.
Figure 2: Numerous applications of RL involve two distinct phases: (1) a training phase, typically unconstrained in time and compute, during which a policy is optimized over a representative distribution of problem instances; and (2) an inference phase, where a limited time and compute budget are allocated to solving a new instance. The inference phase is often overlooked, despite its crucial role in complex tasks where partial observability and the combinatorial growth of observation and action spaces make good solutions unattainable through zero-shot execution alone.
3 Finding the best solution for a given time and compute budget
3.1 Preliminaries
We focus on RL approaches, and use a neural network (policy) that can construct a solution by taking a sequence of actions. This policy is optimised during a training phase and then used during an inference phase, along with an inference strategy, to construct the best possible solution to a new problem instance under a given time and compute budget. These two phases, illustrated in Fig. 2, have different assumptions, objectives and constraints, detailed in the following paragraphs.
Problem instances
We assume that each problem instance can be formulated as a Dec-POMDP (Kaelbling et al., 1998), defined by the tuple $\mathcal{M}=\bigl{(}N,\,\mathcal{S},\,\boldsymbol{\mathcal{O}},\,\boldsymbol{\Omega},\,\boldsymbol{\mathcal{A}},\,R,\,P,\,\gamma,\,H\bigr{)}$. Here $N$ is the number of agents, $\mathcal{S}$ is the environment state space, $\boldsymbol{\mathcal{O}}=\prod_{i=1}^{N}\mathcal{O}^{i}$ is the joint agent observation space, $\boldsymbol{\Omega}:\mathcal{S}\mapsto\boldsymbol{\mathcal{O}}$ is the observation function, $\boldsymbol{\mathcal{A}}=\prod_{i=1}^{N}\mathcal{A}^{i}$ is the joint action space, $R:\mathcal{S}\times\boldsymbol{\mathcal{A}}\mapsto\mathbb{R}$ is the shared reward function, $P:\mathcal{S}\times\boldsymbol{\mathcal{A}}\mapsto\Delta\mathcal{S}$ is the environment transition function, the scalar $\gamma\in[0,1]$ is the discount factor and $H$ is the finite episode horizon. At each timestep the environment occupies a state $s_{t}$ which is mapped to a joint partial observation $\boldsymbol{o}_{t}$ via $\boldsymbol{\Omega}$. The joint action is then sampled following the joint policy, $\boldsymbol{a}_{t}\sim\boldsymbol{\pi}(\,\cdot\mid\boldsymbol{o_{t}})$ and is executed in the environment; after which the environment transitions to the next state $s_{t+1}$, following $P(\,\cdot\mid s_{t},\boldsymbol{a}_{t})$, and the team receives shared the reward $r_{t}=R(s_{t},\boldsymbol{{a}_{t}})$.
Training Phase
We assume a distribution of problem instances $\mathcal{D}$, that can be sampled from. The joint policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}$, parameterised by $\boldsymbol{\theta}$, is used to construct a solution sequentially by taking joint actions conditioned on the joint observation at each timestep. We use RL to train this policy to maximise the expected return obtained when building solutions to instances drawn from the distribution $\mathcal{D}$ over a horizon $H$: $J(\boldsymbol{\pi}_{\boldsymbol{\theta}})=\mathbb{E}_{\mathcal{D}}\left[\sum_{t=0}^{H}\gamma^{t}R(s_{t},\boldsymbol{a_{t}})\right]$. This training objective corresponds to a single attempt (zero-shot). Ideally, this objective should anticipate the multiple attempts allowed at inference, but this is hard to scale. Recent works incorporate such few-shot objectives (Grinsztajn et al., 2023; Chalumeau et al., 2023b, 2024b), but none can yet scale beyond $200$ attempts.
The training phase is usually loosely constrained in terms of time and compute capacity, as typically industrial stakeholders are willing to invest days, weeks or even months of training to obtain a high-performing policy that can generate accurate solutions when deployed in production. Hence, in our experiments, we train all policies until convergence.
Inference Phase
At inference time, a new problem instance $\rho$ is drawn from a distribution $\mathcal{D^{\prime}}$ (possibly different than $\mathcal{D}$). Here, there are typically hard constraints to outputting a final solution: a fixed time limit $\text{T}_{\text{max}}$ constrains wall-clock execution, and a compute capacity $\text{B}_{\text{max}}$ constrains the number of operations that can be done in parallel. The trained policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}$ can be used within these constraints to generate solutions to the problem, and the best solution is ultimately used. The reward function $R$ can still be used to score attempted solutions and inform subsequent attempts. Inference strategies can be defined as a function $\mathcal{I}:(\rho,\boldsymbol{\pi}_{\boldsymbol{\theta}},\text{B}_{\text{max}},\text{T}_{\text{max}})\mapsto(\boldsymbol{a}_{1}^{*},...,\boldsymbol{a}_{H}^{*})$ that uses the base policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}$ and any additional inference-time search, adaptation, storage, or optimisation methods under the budget $\text{T}_{\text{max}}$ and $\text{B}_{\text{max}}$ to produce the best possible solution to the problem instance $\rho$, defined by the sequence of actions $(\boldsymbol{a}_{i}^{*})_{1\leq i\leq H}$. The objective can hence be written as:
$I(\mathcal{I})=\sum_{t=0}^{H}R(s_{t},\boldsymbol{a}_{t}^{*})~{}~{}\text{s.t.}~{}~{}\text{C}(\mathcal{I})\leq\text{B}_{\text{max}},~{}\text{T}(\mathcal{I})\leq\text{T}_{\text{max}}$
where $\text{C}(\mathcal{I})$ and $\text{T}(\mathcal{I})$ represent the compute and time cost of the inference strategy. This formulation highlights that, unlike traditional RL, where zero-shot performance is the primary measure, we focus on strategies which enable further improvement under given constraints.
3.2 One budget, many uses: inference-time search and adaptation
In this section, we detail four types of inference strategies and how we adapt them to work in the multi-agent setting. We implement and release all of these methods in JAX (Bradbury et al., 2018).
Stochastic policy sampling
The first natural lever to improve solution quality is to re-sample from a stochastic policy. In other words, beyond the creation of a unique greedy solution (i.e., using $a=\arg\max_{a^{\prime}}\pi_{\theta}(a^{\prime}|o)$ over a trajectory of observations), one can sample stochastically (as in $a\sim\pi_{\theta}(\cdot|o)$) in order to create diverse solutions. Multi-agent policy sampling generalises easily to the multi-agent case by sampling from the joint action distribution, $\boldsymbol{a}\sim\boldsymbol{\pi}_{\boldsymbol{\theta}}(\cdot|\boldsymbol{o})$.
Tree search
These methods store information about partial solutions using past attempts to preferentially search promising regions of the solution space without updating the pre-trained policy. Simulation guided beam search (sgbs(Choo et al., 2022) provides the best time to performance balance in the literature, outperforming Monte Carlo Tree Search (Coulom, 2006). Like most tree searches, sgbs has three steps: expansion, simulation, pruning. Expansion uses the policy to decide on the most promising next actions (i.e., $)a=\text{top-}K(\pi_{\theta}(\cdot|o))$ from the current node (partial solution). A simulated rollout of an episode is produced greedily using $\pi_{\theta}$ and the return is collected for each node. Pruning keeps only the best nodes found so far based on the return. Solely the expansion step needs to be adapted for Multi-agent SGBS. This is trivial when the explicit joint actions are accessible (de Witt et al., 2020; Yu et al., 2022), since we can still select the top ones (i.e., $\boldsymbol{a}=\text{top-}K(\boldsymbol{\pi}_{\boldsymbol{\theta}}(\cdot|\boldsymbol{o}))$). For methods using auto-regressive action selection (Mahjoub et al., 2025), having access to the top joint actions is intractable, hence we sample K times stochastically from the same node (i.e., $\boldsymbol{a}[1],...,\boldsymbol{a}[K]\sim\boldsymbol{\pi}_{\boldsymbol{\theta}}(\cdot|\boldsymbol{o})$).
Online fine-tuning
These methods keep updating policy parameters at inference time. Given a base policy $\pi_{\theta}$, online fine-tuning optimises $\theta$ using inference-time rollouts and policy gradient updates: $\theta^{\prime}=\theta+\alpha\nabla_{\theta}J(\pi_{\theta})$, where $J(\pi_{\theta})$ represents an adaptation objective. In line with (Bello et al., 2016), we keep maximising expected returns (Bello et al., 2016) over past attempts on the fixed instance (instead of over a training distribution). Multi-agent online fine-tuning re-trains $\boldsymbol{\pi}_{\boldsymbol{\theta}}$ on the new instance using the MARL algorithm that was used during pre-training (Bello et al., 2016; Mahjoub et al., 2025; de Witt et al., 2020; Yu et al., 2022).
Diversity-based approaches
These methods pre-train a collection of diverse specialised policies which can be used to search for the most appropriate solution at inference-time. compass (Chalumeau et al., 2023b) encodes specialised policies in a continuous latent space $\mathcal{L}$ by augmenting a pre-trained policy to condition on both the observation and a latent vector sampled from $\mathcal{L}$ (i.e., $a\sim\pi_{\theta}(\,\cdot\mid o,~{}z),~{}~{}z\sim\mathcal{L}$): effectively creating a continuous collection of policies. compass achieves SOTA in single-agent RL for CO. To avoid having the latent space of Multi-agent COMPASS growing exponentially with the number of agents, we keep one latent space $\mathcal{L}$ for all agents (i.e., $\boldsymbol{a}\sim\boldsymbol{\pi}_{\boldsymbol{\theta}}(\,\cdot\mid\boldsymbol{o},~{}z),~{}~{}z\sim\mathcal{L}$). This allows for tractable training, and for efficient inference search with the covariance matrix adaptation evolution strategy (cma-es(Hansen and Ostermeier, 2001). Aside from being multi-agent, we keep the training and inference phases close to the original method described in Chalumeau et al. (2023b).
4 Experiments
Figure 3: Overview of our tasks and experimental study.
In our experimental study, we combine popular MARL algorithms and inference strategies and benchmark them on a set of complex RL tasks from the literature. Each task was specifically selected for its difficulty. We evaluate all base policies with and without inference-time search across a wide range of budget settings. Our experiments constitute the largest-ever study of inference strategies for decision-making.
Baselines
We use three MARL approaches to obtain our base policies: Independent PPO (de Witt et al., 2020) (ippo) and Multi-Agent PPO (Yu et al., 2022) (mappo), which are widely used and well-known MARL methods, and the recent SOTA sequence modelling approach sable (Mahjoub et al., 2025). Each of these, referred to as base policies, is evaluated with all four inference strategies introduced in Section 3.2, namely stochastic sampling, sgbs, online fine-tuning and compass.
Tasks
Mahjoub et al. (2025) established SOTA over the most comprehensive MARL benchmark published in the field to date. Interestingly, their results demonstrate that there remain certain tasks for which no existing method (including sable) is able to achieve good performance. Specifically, these are tiny-2ag-hard, tiny-4ag-hard, small-4ag, small-4ag-hard, medium-4ag, medium-4ag-hard, medium-6ag, large-4ag, large-4ag-hard, large-8ag, large-8ag-hard, xlarge-4ag and xlarge-4ag-hard from Multi-Robot Warehouse (Papoudakis et al., 2021) (RWARE), smacv2_10_units and smacv2_20_units from the StarCraft Multi-Agent challenge (Samvelyan et al., 2019; Ellis et al., 2023) (SMAC), and con-10x10x10a and con-15x15x23a from Connector (Bonnet et al., 2023).
Figure 4: Training the SOTA algorithm Sable to convergence is not enough to achieve optimal zero-shot performance.
Each environment (illustrated on Fig. 3) introduces distinct challenges that contribute to its complexity. RWARE requires agents to coordinate in order to pick up and deliver packages without collision and has a very sparse reward signal. In SMACv2 tasks, a team cooperates in real-time combat against enemies across diverse scenarios with randomised generation. Connector models the routing of a printed circuit board where agents must connect to designated targets without crossing paths. All three environments feature combinatorial and high-dimensional action spaces, partial observability and the need for tightly coordinated behaviours, making these 17 tasks a compelling test-bed for complex RL with desirable properties modelling aspects of real-world tasks. We use JAX-based implementations of Connector and RWARE from Jumanji (Bonnet et al., 2023) and for SMAC from JaxMARL (Rutherford et al., 2023).
Training base policies
To obtain clear performance ceilings for each algorithm and best isolate the effects of inference strategies, we train all base policies until convergence. For the sake of continuity with previous work with truncated training budgets, typically of $20$M steps, we report the zero-shot results for each converged checkpoint compared to its previous corresponding reported performance on Fig. 4. We observe that in most tasks (14 out of 17), the converged policy stays below 70% normalised performance. compass requires an additional training phase, which reincarnates the existing base policies to create the latent space specialisation. For each base policy and task, we also train the compass checkpoint until convergence. This leads to $102$ trained policy checkpoints.
Evaluating performance during inference
Evaluating inference strategies in a way that is unbiased and aligned with real-world settings is challenging. Most papers report results where the budget is based on a number of attempts, hence not directly incorporating the time cost of the strategy. The time costs are reported, but it is tough to analyse due to the plurality of hardware used to obtain them. Having re-implemented all of the baselines in the same code base and setting the budget in terms of time (in seconds), we can avoid this bias in our study. We use the same fixed hardware for all our experiments, namely a NVIDIA-A100-SXM4-80GB GPU. For statistical robustness, we always run $128$ independent seeds. In all cases, we control for $\text{B}_{\text{max}}$ by varying the permitted number of batched parallel attempts instead of altering hardware between experiments. For aggregation across multiple tasks we follow the recommendations made by Agarwal et al. (2021) and use the rliable library to compute and report the inter-quartile mean (IQM) and 95% stratified bootstrap confidence intervals.
Hyperparameters
To train the base policies, we re-use the hyperparameters reported in Mahjoub et al. (2025), which have been optimised for our tasks. For the inference strategies, we follow recommendations from the literature (Choo et al., 2022; Chalumeau et al., 2023b). All hyperparameters choices are reported in Appendix C.
4.1 A couple of seconds is all you need
In this section, we demonstrate that inference-time search can help reach close to maximum task performance, using base policies for which zero-shot performance stagnates around $60$%.
Figure 5: Performance obtained by inference strategies over the benchmark. Each base policy is evaluated with each possible inference strategy. We report the inter-quartile mean over tasks with 95% stratified bootstrap confidence intervals.
Experiments
To demonstrate that inference-time strategies are accessible, we use a small budget: $30$ seconds, and a compute capacity enabling to generate $64$ solutions in parallel. Each base policy is evaluated greedily for a single attempt, and then evaluated with the search budget, using each inference strategy. We report the performance distribution over the $17$ tasks in Fig. 5, and the performance gains offered by the best inference-time search over the best zero-shot on Fig. 1.
Discussion
We can draw four main conclusions. First, inference-time search does provide a massive performance boost over zero-shot, which stands for every base policy. For the SOTA zero-shot method sable, this translates to pushing the best-ever achieved aggregated performance by more than $45$% and creating a system (sable +compass) that achieves close to $100$% win-rate in all tasks where this metric is available. Second, the improvement enabled over zero-shot performance increases significantly (almost exponentially) with respect to the complexity of the task (see Fig. 1). This suggests substantial gains are still ahead as the field moves toward increasingly realistic scenarios. Third, we observe that compass is the leading strategy across tasks and base algorithms, and that sable remains the SOTA base policy even when using inference-time search. Interestingly, under a small time budget, stochastic sampling outperforms tree search and online fine-tuning. We nevertheless show in the following section that this order can be reversed when more budget is provided.
4.2 Mapping performance with compute and time budget
A recurrent limitation in previous work on inference strategies is the use of a fixed budget during evaluation, creating a narrow view on methods, and often creating a bias towards certain types of methods. In this section, we aim at providing a much broader perspective over inference-time search by reporting performance over a grid of time and compute budgets.
Experiments
We choose a maximum time of $300$ seconds, and evaluate all inference strategies using the leading base policy (sable), with a compute budget of $\{4,8,16,32,64,128,256\}$. All in all, we have $4$ inference strategies, $7$ possible compute budgets, $17$ tasks, and $128$ seeds per task, leading to 60 9286092860\,92860 928 evaluated episodes. This constitutes the largest study released on inference strategies. We report these results using contour plots, where the $x$-axis is time, $y$-axis the number of parallel attempts allowed (our proxy for compute) and colour corresponds to the performance achieved (win-rate when accessible or min-max normalised return) going from dark purple (min) to yellow (max). We keep the $8$ hardest tasks, the lower half based on the zero-shot performance of the converged sable checkpoints (see Fig. 4), on Fig. 6 and defer remaining tasks to Appendix A.
Figure 6: Performance contour plots of Sable with different inference strategies on a range of time budget ($x$-axis) and compute capacities ($y$-axis), for a set of representative tasks. Colours indicates performance, with brighter colours indicating higher values.
Discussion
As a sanity check, we remark that performance always increases (colours become lighter) when time or compute increases (going towards the upper right corner). We now highlight three main observations. First, compass demonstrates impressive versatility and achieves significant gains over other inference strategies, dominating all maps, except for con-10x10x10a, where it gets slightly outperformed by online fine-tuning. Second, we observe high variance for online fine-tuning: getting close to compass for large budgets on con-10x10x10a, yet struggling to match stochastic sampling on others (e.g., RWARE’s large-8ag). This shows that fine-tuning can be detrimental by reducing the number of attempts made within the time budget. Plus, policy gradients can be unstable (small batch size) or converge to local optima. This observation disproves the common belief that inference-time search is as trivial as over-fitting to the problem instance. Finally, we observe that sgbs has good performance for high compute capacity and low time, but the greediness of its simulation step impacts its exploration capacity, preventing to benefit from additional time budget.
4.3 Scaling with increasing budget
In practical applications, time is often more restricted than compute: a couple of seconds or minutes can be allowed, sometimes a few hours (train scheduling for the next day), but rarely more. Being able to improve solutions’ quality under a fixed time budget by scaling compute is a valuable property. Consequently, methods that can scale with the compute available are particularly valuable. We look at this property in this section.
Experiments
We keep the time budget fixed, $300$ seconds, and we plot the final performance for each possible compute budget (still using the number of parallel attempts as a proxy). We use sable as the base policy and evaluate all the inference strategies across the $8$ hardest tasks. We report results per strategy, and aggregated over the tasks on Fig. 7.
Figure 7: Scaling of performance with respect to compute capacity, for a fixed time budget. Compute capacity ($x$-axis) is set by limiting the number of parallel attempts. sable +compass demonstrates impressive scaling properties, reaching the performance bounds of the benchmark.
Discussion
As expected, stochastic sampling has the lowest scaling coefficient: its search solely relies on chance, no adaptation or additional search is happening. On the other hand, we can see that online fine-tuning benefits from more compute, probably due to a better estimation of the policy gradient. It nevertheless requires a budget of 64 parallel attempts to clearly outperform stochastic sampling. On the other hand, compass consistently provides a significant advantage. Its scaling trend seems linear at first, with a high coefficient, and only seems to decline for higher budgets, as performance limits are reached (over $95$% win-rate when accessible, best-ever observed performance elsewhere). We have a two-fold explanation for these particularly impressive scaling properties: (i) the diversity contained in the latent space can be exploited by more parallelism, leading to a massive exploration of the solution space, even when the initial policy is far from optimal, and (ii) the higher the batch size, the better the cma-es search, enabling compass to exploit even more information from any additional searching step allowed within the given time budget.
5 Conclusion
In this work, we demonstrate that inference-time strategies are a critical and underutilized lever for boosting the performance of RL systems in complex tasks, using multi-agent RL as a representative test-bed. While training-time improvements have long dominated the field, our results show that inference-time search may offer significant performance gains using only a few seconds to minutes of additional wall-clock time during execution. We introduce a unified framework for inference strategies, extend it to the multi-agent setting, and empirically validate its effectiveness under varying compute and time budgets. Our large-scale evaluation, the most comprehensive to date on inference-time methods, reveals three key takeaways: (i) inference-time search with a relevant strategy yields significant improvements, even under tight time constraints; (ii) the gains depend on the inference budget, and our contour maps provide practitioners with practical guidance based on their constraints; and (iii) sable +compass not only dominates the benchmark but also exhibits the most favourable scaling trends, making it particularly effective for increasingly complex decision-making problems.
Altogether, our findings call for a shift in how decision-making models are evaluated and deployed: inference-time strategies should be treated as core components of the solution pipeline, not as optional refinements. We hope our results and open-source tools will encourage broader adoption and inspire further innovation in the design of scalable inference-time algorithms.
Limitations and future work
We focus on multi-agent RL to represent complex RL. We are aware that single-agent tasks can be arbitrarily complex. We believe that our main trends should generalise to the single-agent case, but leave it for future work. We intend two main future research directions. First, studying how to best combine existing paradigms together. Second, investigating how inference strategies reacts to evaluation out-of-distribution.
Acknowledgements
We would like to thank Guillaume Toujas-Bernate, Jake Lourie and Thomas Lecat for useful discussions on the use of inference strategies in real-world applications. We thank our MLOps team for developing our model training and experiment orchestration platform AIchor. We thank the Python and JAX communities for developing tools that made this research possible. Finally, we thank Google’s TPU Research Cloud (TRC) for supporting our research with Cloud TPUs.
Appendix
Appendix A Additional results
Section 4.2 presents the contour plots of Sable with all inference strategies on the hardest 8 tasks of the benchmark. We report the contour plots of the other $9$ remaining tasks on Fig. 8.
Figure 8: Performance contour plots of Sable with different inference strategies on remaining 9 tasks on a range of time budget ($x$-axis) and compute capacities ($y$-axis). Colours indicates performance, with brighter colours indicating higher values.
Fig. 8, shows similar trends to those illustrated in Section 4.2. sable +compass leads to the best overall performance. Stochastic sampling provides a good and robust baseline overall. Finally, sgbs and online fine-tuning suffer from variance: despite being able to compete close to compass on a couple tasks, they often get outperformed by stochastic sampling. We do not report the results for sable +sgbs on smacv2_10_units because we found an issue, but the figure will be updated shortly with new results.
Appendix B Details about the tasks
Figure 9: Environment rendering for Connector. Task name: con-10x10-10a. Image from Mahjoub et al. [2025].
Connector
is an environment designed to model the routing of a printed circuit board (PCB). Agents start at randomly generated initial positions and have to reach a goal position without overlapping with each other’s paths (indicated as lower opacity coloured cells in Figure 9). Each agent has a partial view with a fixed field of view around itself as well as its current $(x,y)$-coordinate on the grid and the $(x,y)$-coordinate of its goal location. At each timestep, agents receive a small negative reward $-0.03$ and receive a reward of $1$ for successfully connecting to a goal location. The particular difficulty in this environment stems from the fact that agents have to select actions carefully so that they not only reach their goal greedily but so that all agents can ultimately reach their goals.
Figure 10: Environment rendering for Robot Warehouse. Task name: tiny-2ag. Image from Mahjoub et al. [2025].
RWARE
is an environment where agents need to coordinate to deliver shelves (green cells in Figure 10) to a goal location (dark grey cells in Figure 10) in a warehouse. The reward is sparse since agents only receiving a reward of $1$ for a successful shelf delivery to the goal location and 00 otherwise. This sparsity makes the environment particularly difficult since agents have to complete a long sequence of correct actions in order to receive any reward signal.
Figure 11: Environment rendering for SMAX. Task name: 2s3z. Image from Mahjoub et al. [2025].
SMAC
is an environment where agents need to coordinate and develop the correct micro strategies in order to defeat an enemy team. It was shown that the original benchmark was overly simplistic and for this reason SMACv2 [Ellis et al., 2023] was developed in order to address the shortcomings of the original benchmark. SMACv2 randomly generates a team of allied units by sampling from a selection of unit types and randomly generates ally starting positions at the start of each episode. The need for generalisation is what gives this benchmark its difficulty.
JAX-based implementations.
Since Connector was developed as part of Jumanji [Bonnet et al., 2023], we make use of the JAX-based implementation directly. For RWARE we use the JAX-based reimplementation from Jumanji and for SMAC and SMACv2 we make use of the JAX-based reimplementation from JaxMARL [Rutherford et al., 2023], named SMAX. For a detailed discussion on task naming conventions we refer the reader to the Appendix of Mahjoub et al. [2025].
Appendix C Hyper-parameters
In this section, we report the hyper-parameters used in our experiments. They can also be found in the submitted code files.
Training - base policies
Hyper-parameters used to train the base policies for this benchmark are taken from Mahjoub et al. [2025]. These hyper-parameters were tuned on each task with a budget of 40 trials using the Tree-structured Parzen Estimator (TPE) Bayesian optimisation algorithm, and are reported in Mahjoub et al. [2025].
Hyper-parameters - COMPASS Training
Hyper-parameters used to train all the compass checkpoints can be found in Table 1. Most of them were kept close to the original hyper-parameters used in Chalumeau et al. [2023b].
Table 1: Hyper-parameters used for compass’ training phase. The same values are used for all three base policies (i.e., sable, ippo, mappo).
Parameter Value
Instances batch size 128
Latent space sample size 64
Latent space dimension size 16
Latent amplifier 1
Padding with random weights True
Weight noise amplitude 0.01
Hyper-parameters - Inference-time search
Most values are directly taken from original works or kept close to these. When different, we tried to use guidance from original work to decide how to set them. The values for the cma-es search used with compass, for all base policies, can be found in Table 2. The values for online fine-tuning (all base policies) are reported in Table 3. Hyper-parameters for sgbs (all base policies) are reported in Table 4, and for stochastic sampling in Table 5.
Table 2: Hyper-parameters of the cma-es process used to search compass’ latent space at inference time. The same parameters are used for all tasks and all base policies (i.e., sable, mappo and ippo).
Parameter Number of attempts
4 8 16 32 64 128 256
Latent sample size 4 8 16 32 64 128 256
Number of elites 2 4 8 16 32 64 128
Covariance matrix step size 1.0 1.0 1.0 1.0 1.0 1.0 1.0
Initial sigma 1.0 1.0 1.0 1.0 1.0 1.0 1.0
Num. components 1 1 1 1 1 1 1
Table 3: Hyper-parameters used when doing online fine-tuning at inference-time with sable, mappo and ippo.
Parameter Task Name Sable IPPO MAPPO
Max. trajectory size All tasks 64 64 64
Learning rate con-10x10x10a 0.001 0.00025 0.0001
con-15x15x23a 0.001 0.0001 0.0001
large-4ag 0.001 0.0005 0.0001
large-4ag-hard 0.001 0.0005 0.00025
large-8ag 0.001 0.00025 0.0001
large-8ag-hard 0.001 0.00025 0.0005
medium-4ag 0.001 0.00025 0.00025
medium-4ag-hard 0.0005 0.0005 0.00025
medium-6ag 0.0005 0.0001 0.00025
smacv2_10_units 0.00025 0.0005 0.00025
smacv2_20_units 0.001 0.0005 0.0005
smacv2_5_units 0.001 0.00025 0.00025
small-4ag 0.0005 0.0005 0.0005
small-4ag-hard 0.001 0.0001 0.00025
tiny-2ag-hard 0.001 0.00025 0.00025
tiny-4ag-hard 0.0005 0.0005 0.0005
xlarge-4ag 0.001 0.0005 0.0005
xlarge-4ag-hard 0.0001 0.0005 0.0001
Entropy coefficient Connector 0.05 0.00 0.00
RWARE 0.0 0.0 0.0
SMAX 0.0 0.0 0.0
Table 4: Hyper-parameters for Simulation Guided Beam Search (sgbs). The same values are used for sable, ippo and mappo. Values are chosen such that beam width and top-k use fully the number of attempts allowed by batch.
Parameter Number of attempts
4 16 32 64 128 256
Beam width 4 4 4 4 4 4
Top k 1 4 8 16 32 64
Table 5: To sample stochastically from a policy (sable, ippo or mappo), we always use the same temperature as the one used during training: $1.0$.
Parameter Value
Evaluation greedy False
Temperature 1.0
Appendix D Experimental stack
All algorithmic implementations are built by extending the JAX-based research library, Mava [de Kock et al., 2021]. Our version of cma-es used for searching compass’s latent space comes from the JAX-based quality diversity library, QDAX [Chalumeau et al., 2024a]. Both these libraries are built to leverage the DeepMind JAX ecosystem [DeepMind et al., 2020] and use Flax [Heek et al., 2024] for building neural networks, optax for optimisers and orbax for model checkpointing.
Appendix E Additional details about training
E.1 Base policy training
In Section 3, we explain that in our setting, it is assumed that all methods can be trained until convergence. For transparency and comparison with previous results in the literature, we report in  Table 6 the training budget given to the methods to reach convergence, in environment steps.
Table 6: Training budgets (in environment steps) for each method across tasks.
Task Name Sable MAPPO IPPO
con-10x10x10a 100M 200M 200M
con-15x15x23a 100M 200M 200M
large-4ag 600M 200M 200M
large-4ag-hard 600M 400M 200M
large-8ag 600M 400M 200M
large-8ag-hard 600M 200M 200M
medium-4ag 400M 200M 200M
medium-4ag-hard 600M 200M 200M
medium-6ag 600M 200M 200M
small-4ag 200M 200M 200M
small-4ag-hard 400M 200M 200M
tiny-2ag-hard 100M 200M 200M
tiny-4ag-hard 200M 200M 200M
xlarge-4ag 200M 400M 200M
xlarge-4ag-hard 200M 200M 200M
smacv2_10_units 100M 200M 200M
smacv2_20_units 600M 200M 200M
E.2 COMPASS training
In this section, we provide details about compass and our implementation. For a more in-depth explanation, we refer the reader to the original paper [Chalumeau et al., 2023b]. Our implementation is available in the submitted code files.
compass is a method, which (i) uses a pre-trained base policy and modifies it to be conditioned by a latent vector (ii) re-trains it to enable latent vectors to create policies that are diversified and specialised for distinct training sub-distributions. At inference-time, this enables to search for the latent vector which performs best on the new instance to be solved.
Reincarnating a pre-trained base policy
compass checkpoints are not trained from scratch, they reincarnate an existing base policy, add parameters to this policy in order to process the latent vector which can be given as additional input. Hence, one must choose where to inject the latent vector in the existing architecture, and must modify the existing parameters of the neural policy to account for the new shape of the input. The latent vector must be input in a way that enables to diversify the actions produced for the same observation (i.e., making sure that different latent vectors create different policies). In practice, we concatenate the latent vector to the observations. When relevant, we make sure that this is done after observation normalisation layers.
In the original work, the additional weights added to the base neural network are initialised with zeros. In our case, we observed that this could not provide enough diversification in the first training step, preventing any emergence of specialisation, leading to no benefit. We found that initialising these parameters with random uniform values between $-0.01$ and $0.01$ fixed the issue across all tasks.
Creating diversity and specialisation in the latent space
Similarly to the original method, the latent space is not learned, it is a fixed prior, and the network’s weights are learned to create diversity from this prior space. We also use a similar prior, which is a uniform distribution between $-1$ and $1$, over $16$ dimensions. We do not need any amplification of the latent vector (i.e., we multiply by $1$, whereas the original work multiplies it by $100$). At each training step, a batch of instances is sampled from the training distribution, a batch of latent vectors is sampled from that latent space, the conditioned policy is evaluated for each instance, for each latent vector, and only the best performing latent vectors, for each instance, are used in the computation of the loss. This creates the specialisation (diversification). For additional motivation and mathematical formulation, we refer the reader to Chalumeau et al. [2023b]. Like all training processes in our experimental study, the method is allowed to train until convergence. In practice, each compass checkpoint was trained for $100$ million steps.
Appendix F Pseudo-algorithms
In this section, we provide algorithmic descriptions for the inference strategies used in the paper, showing how the policy is used at inference time, under the time and budget constraints. Stochastic sampling is explained in Algorithm 1, sgbs in Algorithm 2, online fine-tuning in Algorithm 3 and compass in Algorithm 4.
Algorithms use the terminology trajectory $\tau$, which is equivalent to solution. The symbol $\circ$ is used for the concatenation of a new step transition to a partial trajectory.
Algorithm 1 Stochastic Sampling
1:  Input: policy $\boldsymbol{\pi}_{\theta}$, instance $\rho$, time budget $T_{\max}$, compute capacity $B_{\max}$, reward function $\mathcal{R}$
2:  Initialize best trajectory $\tau^{*}\leftarrow\emptyset$, best score $R^{*}\leftarrow-\infty$
3:  while elapsed time $<T_{\max}$ do
4:     for parallel rollout $b=1$ to $B_{\max}$ do
5:        Initialize trajectory $\tau_{b}\leftarrow\emptyset$
6:        for step $t=1$ to $H$ do
7:           Observe $\boldsymbol{o}_{t}$ from environment copy $\rho_{b}$
8:           Sample $\boldsymbol{a}_{t}\sim\boldsymbol{\pi}_{\theta}(\cdot\mid\boldsymbol{o}_{t})$
9:           Execute $\boldsymbol{a}_{t}$ in $\rho_{b}$, observe $\boldsymbol{o}_{t+1}$
10:           Append $(\boldsymbol{o}_{t},\boldsymbol{a}_{t})$ to $\tau_{b}$
11:        $R_{b}\leftarrow\mathcal{R}(\tau_{b})$
12:        if $R_{b}>R^{*}$ then
13:           $\tau^{*}\leftarrow\tau_{b}$,  $R^{*}\leftarrow R_{b}$
14:  return $\tau^{*}$
Algorithm 2 Simulation-Guided Beam Search
1:  Input: policy $\boldsymbol{\pi}_{\theta}$, instance $\rho$, time budget $T_{\max}$, compute capacity $B_{\max}$, beam width $K$, reward function $\mathcal{R}$
2:  Initialize beam ${}\mathcal{B}\leftarrow\{\emptyset\}$, best trajectory $\tau^{*}\leftarrow\emptyset$, best score $R^{*}\leftarrow-\infty$
3:  while elapsed time $<T_{\max}$ do
4:     $\mathcal{B}_{\text{new}}\leftarrow\emptyset$
5:     for partial trajectory $\tau$ in $\mathcal{B}$ do
6:        for sample $b=1$ to $B_{\max}/|\mathcal{B}|$ do
7:           Let $\boldsymbol{o}_{t}$ be the observation at the end of $\tau$
8:           Sample $\boldsymbol{a}_{t}\sim\boldsymbol{\pi}_{\theta}(\cdot\mid\boldsymbol{o}_{t})$
9:           Simulate greedy rollout from $\tau\circ\boldsymbol{a}_{t}$ using $\boldsymbol{\pi}_{\theta}$
10:           Let $\hat{\tau}_{b}$ be the resulting full trajectory
11:           $R_{b}\leftarrow\mathcal{R}(\hat{\tau}_{b})$
12:           Add $(\tau\circ\boldsymbol{a}_{t},R_{b})$ to $\mathcal{B}_{\text{new}}$
13:           if $R_{b}>R^{*}$ then
14:              $\tau^{*}\leftarrow\hat{\tau}_{b}$,  $R^{*}\leftarrow R_{b}$
15:     Prune top-$K$ trajectories from $\mathcal{B}_{\text{new}}$ into $\mathcal{B}$
16:  return $\tau^{*}$
Algorithm 3 Online Fine-Tuning
1:  Input: base policy $\boldsymbol{\pi}_{\theta}$, instance $\rho$, time budget $T_{\max}$, compute capacity $B_{\max}$, reward function $\mathcal{R}$, learning rate $\alpha$
2:  Initialize best trajectory $\tau^{*}\leftarrow\emptyset$, best score $R^{*}\leftarrow-\infty$
3:  Initialize adapted parameters $\theta^{\prime}\leftarrow\theta$
4:  while elapsed time $<T_{\max}$ do
5:     for parallel rollout $b=1$ to $B_{\max}$ do
6:        Rollout $\tau_{b}\sim\boldsymbol{\pi}_{\theta^{\prime}}$ on $\rho$
7:        $R_{b}\leftarrow\mathcal{R}(\tau_{b})$
8:        if $R_{b}>R^{*}$ then
9:           $\tau^{*}\leftarrow\tau_{b}$,  $R^{*}\leftarrow R_{b}$
10:     Compute gradient: $\nabla_{\theta}\leftarrow\frac{1}{B_{\max}}\sum_{b=1}^{B_{\max}}\nabla_{\theta}\log\boldsymbol{\pi}_{\theta^{\prime}}(\tau_{b})\cdot R_{b}$
11:     Update parameters: $\theta^{\prime}\leftarrow\theta^{\prime}+\alpha\nabla_{\theta}$
12:  return $\tau^{*}$
Algorithm 4 COMPASS + CMA-ES search
1:  Input: latent-conditioned policy $\boldsymbol{\pi}_{\theta}(\cdot\mid\boldsymbol{o},z)$, instance $\rho$, time budget $T_{\max}$, compute capacity $B_{\max}$, reward function $\mathcal{R}$
2:  Initialize CMA-ES: mean $\mu$, covariance $\Sigma$
3:  Initialize best trajectory $\tau^{*}\leftarrow\emptyset$, best score $R^{*}\leftarrow-\infty$
4:  while elapsed time $<T_{\max}$ do
5:     Sample $\{z_{b}\}_{b=1}^{B_{\max}}\sim\mathcal{N}(\mu,\Sigma)$
6:     for each $z_{b}$ in parallel do
7:        Rollout $\tau_{b}\sim\boldsymbol{\pi}_{\theta}(\cdot\mid\boldsymbol{o},z_{b})$ on $\rho$
8:        $R_{b}\leftarrow\mathcal{R}(\tau_{b})$
9:        if $R_{b}>R^{*}$ then
10:           $\tau^{*}\leftarrow\tau_{b}$,  $R^{*}\leftarrow R_{b}$
11:     Update $(\mu,\Sigma)$ using CMA-ES with $\{(z_{b},R_{b})\}_{b=1}^{B_{\max}}$
12:  return $\tau^{*}$
Appendix G Discussion about the inference-time budget
Choice of the budget used in the experiments
We create our experimental setting to be similar to practical use cases. Hence, we use time as the main budget constraint, instead of number of attempts (contrary to most literature on the topic). This enables to account for the cost of search and adaptation. Even though we always evaluate methods on $128$ instances for each task, we ensure that each instance is solved fully independently, to avoid that the vectorisation process creates a bias towards one method or another. We choose to study methods with time budgets ranging from $30$ to $300$ seconds. The lower bound is taken to be small, to show that even short inference-time searches can provide significant benefit. The higher bound is chosen to be high enough to see which method can really benefit from a longer search budget. Concerning the compute capacity, we choose a range of values that seems achievable in practice. In real-world scenarios, simulations can get more complex and require one CPU core for one simulation. We hence believed that most use cases would fit within the limit of $256$ simulations in parallel. Nevertheless, our code works for higher values and can be used to obtain results in new settings (both in terms of time and parallel attempts).
Total number of attempts achieved within time budget
To remain consistent with previous literature, and provide a better understanding of the search and adaptation cost of each inference strategy and each base policy, we provide the number of attempts achieved within the time limit, for all tasks and compute capacity (i.e., number of parallel attempts allowed). Values for sable within 300 seconds are reported on Table 7 for compass, Table 8 for online fine-tuning, Table 9 for stochastic sampling. sgbs builds new solutions from existing partial solutions, hence the count of the attempted solution cannot be obtained in the same way as other methods.
Values for all base policies within 30 seconds are reported on Table 10Table 12 and Table 11.
Table 7: Number of attempts (solutions generated) when using sable + compass during 300 seconds of inference-time search.
Task Name Number of attempts
4 8 16 32 64 128 256
con-10x10x10a 9960 19448 37520 67296 131456 248320 422144
con-15x15x23a 1608 3112 7744 14144 27456 53248 80896
large-4ag 2608 5208 9360 16896 30464 44032 59392
large-4ag-hard 2596 5000 9360 16832 29568 44288 59904
large-8ag 1612 3200 6240 10304 17280 23424 30976
large-8ag-hard 1708 3144 6096 10624 18880 29056 39680
medium-4ag 2088 4072 7904 15104 29376 48256 87296
medium-4ag-hard 2072 4040 7888 16992 29632 48384 87552
medium-6ag 1536 3016 5728 10816 21440 37760 74240
smacv2_10_units 7628 13248 26176 45088 74304 122624 174336
smacv2_20_units 5824 10428 18528 34112 64064 111232 180992
small-4ag 1964 3880 8672 14112 28096 61568 83456
small-4ag-hard 1876 4216 7104 15488 30464 52096 93696
tiny-2ag-hard 2736 6072 11856 22496 44544 77440 147200
tiny-4ag-hard 1384 2648 5024 10624 16640 32256 45568
xlarge-4ag 1816 3568 7952 15264 24320 45440 81152
xlarge-4ag-hard 1180 2264 5344 9696 17792 28544 51200
Table 8: Number of attempts (solutions generated) when using online fine-tuning and sable during 300 seconds of inference-time search.
Task Name Number of attempts
4 8 16 32 64 128 256
con-10x10x10a 10608 20256 37744 64448 115392 226560 395776
con-15x15x23a 2136 3952 6864 10496 16592 32000 60928
large-4ag 1792 3296 6560 9920 13376 24448 49664
large-4ag-hard 2064 3416 6448 8832 13248 24448 46592
large-8ag 1416 2336 4592 7552 11136 19200 34304
large-8ag-hard 1140 2576 4064 6752 12736 26624 44032
medium-4ag 2000 4176 7472 9920 15744 28160 48384
medium-4ag-hard 2024 3776 7232 10816 14848 28160 48384
medium-6ag 1700 2872 6448 8704 21120 32896 65792
smacv2_10_units 6848 12576 20176 37536 68160 156672 235776
smacv2_20_units 5700 10584 18176 25216 47488 104382 155904
small-4ag 1940 3688 7072 9600 14272 30336 49408
small-4ag-hard 1940 3688 7072 9600 14272 30336 49408
tiny-2ag-hard 2572 5600 9824 13344 22848 48768 86272
tiny-4ag-hard 2572 5568 9808 13344 22784 48768 86272
xlarge-4ag 1768 3856 6368 9792 13760 24448 46080
xlarge-4ag-hard 1280 2024 3776 6368 11456 21504 40192
Table 9: Number of attempts (solutions generated) when sampling stochastically from sable during 300 seconds of inference-time search.
Task Name Number of attempts
4 8 16 32 64 128 256
con-10x10x10a 10632 20504 39524 70720 138624 264960 448512
con-15x15x23a 2048 4056 7840 14144 27648 54144 101376
large-4ag 2092 3624 7808 15456 30336 46464 82944
large-4ag-hard 2108 3696 8112 13344 26816 46592 82176
large-8ag 1508 2552 5824 9568 17984 30720 52480
large-8ag-hard 1324 2256 4432 8416 16576 34048 32208
medium-4ag 2388 4576 8064 15296 29760 53120 81664
medium-4ag-hard 2144 4616 7968 16992 29696 52992 88576
medium-6ag 1344 3496 6768 10848 25280 66048 66048
smacv2_10_units 7168 11056 25376 47088 94720 148608 261888
smacv2_20_units 4848 9144 17584 34688 55232 102144 228608
small-4ag 1472 3896 7072 16448 27840 61056 84224
small-4ag-hard 1812 3704 7072 16448 26880 59776 84224
tiny-2ag-hard 2824 6168 10432 20192 45504 77952 165632
tiny-4ag-hard 1600 2672 5888 9184 16960 33152 58880
xlarge-4ag 1832 3624 7888 13248 27584 50944 81152
xlarge-4ag-hard 1164 2280 4480 8064 15296 33152 58880
Table 10: Number of attempts (solutions generated) when using compass with sable, ippo and mappo during 30 seconds of inference-time search.
Task Name Number of attempts - Sable PPO
4 8 16 32 64 128 256 64
con-10x10x10a 996 1872 3744 6720 13120 24832 42240 65152
con-15x15x23a 160 384 768 1408 2752 5376 8192 36608
large-4ag 260 464 928 1664 3072 4480 5888 5056
large-4ag-hard 256 464 928 1664 2944 4480 6144 4928
large-8ag 160 312 624 1024 1728 2304 3072 3520
large-8ag-hard 168 304 608 1056 1856 2944 4096 3584
medium-4ag 208 392 784 1504 2944 4864 8704 5120
medium-4ag-hard 208 392 784 1696 2944 4864 8704 5056
medium-6ag 152 288 576 1088 2112 3840 7424 4416
smacv2_10_units 760 1304 2608 4480 7424 12288 17408 30464
smacv2_20_units 584 928 1856 3392 6400 11136 18176 26880
small-4ag 196 432 864 1408 2816 6144 8320 5248
small-4ag-hard 188 352 704 1536 3072 5120 9216 5248
tiny-2ag-hard 272 592 1184 2240 4480 7680 14592 6848
tiny-4ag-hard 136 248 496 1056 1664 3200 4608 5312
xlarge-4ag 180 392 784 1536 2432 4480 8192 4160
xlarge-4ag-hard 116 264 528 960 1792 2816 5120 4288
Table 11: Number of attempts (solutions generated) when using stochastic sampling with sable, ippo and mappo during 30 seconds of inference-time search.
Task Name Number of attempts - Sable PPO
4 8 16 32 64 128 256 64
con-10x10x10a 1060 2048 3940 7040 13760 26240 44800 103680
con-15x15x23a 204 400 780 1408 2752 5376 10112 51456
large-4ag 208 360 768 1536 3008 4608 8192 5312
large-4ag-hard 208 368 800 1312 2688 4608 8192 5184
large-8ag 148 248 576 928 1792 3072 5120 3904
large-8ag-hard 132 224 432 832 1664 3328 3200 3712
medium-4ag 236 456 800 1504 2944 5248 8128 5120
medium-4ag-hard 212 456 784 1696 2944 5248 8832 5312
medium-6ag 132 344 672 1056 2496 6528 6528 4480
smacv2_10_units 716 1104 2528 4704 9408 14848 26112 33408
smacv2_20_units 484 912 1744 3456 5504 10176 22784 31808
small-4ag 144 384 704 1632 2752 6016 8384 5184
small-4ag-hard 180 368 704 1632 2688 5888 8384 5248
tiny-2ag-hard 280 608 1040 2016 4544 7808 16512 4544
tiny-4ag-hard 160 264 576 896 1664 3328 5888 5312
xlarge-4ag 180 360 784 1312 2752 5120 8128 4416
xlarge-4ag-hard 116 224 448 800 1536 3328 5888 4352
Table 12: Number of attempts (solutions generated) when using online fine-tuning with sable, ippo and mappo during 30 seconds of inference-time search.
Task Name Number of attempts - Sable PPO
4 8 16 32 64 128 256 64
con-10x10x10a 1060 2024 3760 6432 11520 22656 39424 23680
con-15x15x23a 212 392 672 1024 1600 3200 5888 5888
large-4ag 176 328 656 992 1280 2432 4864 3136
large-4ag-hard 204 336 640 864 1280 2432 4608 2944
large-8ag 140 232 448 736 1088 1920 3328 2304
large-8ag-hard 112 256 400 672 1216 2560 4352 2112
medium-4ag 200 416 736 992 1536 2816 4608 2880
medium-4ag-hard 200 376 720 1056 1472 2816 4608 2880
medium-6ag 168 280 640 864 2112 3200 6400 4160
smacv2_10_units 684 1256 2016 3744 6784 15616 23552 8960
smacv2_20_units 568 1056 1808 2496 4736 10368 15360 7936
small-4ag 192 368 704 960 1408 2944 4864 2816
small-4ag-hard 192 368 704 960 1408 2944 4864 2816
tiny-2ag-hard 256 560 976 1312 2240 4864 8448 3904
tiny-4ag-hard 256 552 976 1312 2240 4864 8448 3904
xlarge-4ag 176 384 624 960 1344 2432 4608 2944
xlarge-4ag-hard 128 200 368 608 1088 2048 3840 2432